Can you please help me solve this equation, i have gotten this far is this correct ?
5t/12=4-3t
5t/12*12=4-3t*12
5t=48+36t
5t+36=48
41t=48
41t/41=48/41
t= 48/41
t=1.17
B: Prove your solution is correct?
thank you :)

Found 2 solutions by Maths68, ptaylor:
Answer by Maths68(1474)

5t/12=4-3t
5t/12*12=4-3t*12
5t=48-36t ---------------> here you made a little mistake you changed sign (- to +)
5t+36=48
41t=48
41t/41=48/41
t= 48/41
t=1.17


You did all, Well Done!


I will only check your answer is right or wrong


Check
=======
5t/12=4-3t
put the value of t=48/41 (that is your answer in fraction)
5(48/41)/12=4-3(48/41)
Multiply by 12 as you did earlier
5(48/41)=48-36*(48/41)
now multiply by 41 both sides
41*(5(48/41)=41*(48-36*(48/41))
5(48)=41*48-(36*48)
240=1968-1728
240=240

===========================================
5t/12=4-3t
put the value of t=1.17 (that is your answer in decimal)
5(1.17)/12=4-3(1.17)
5.85/12=4-3.51
0.49=0.49

So you found the correct value of t which is 48/41 or 1.17


Answer by ptaylor(2198)

YOU DID GREAT!!! COUPLE OF MISTAKES BUT YOU RECOVERED.
5t/12=4-3t
5t/12*12=4-3t*12
5t=48+36t--------------------shud be -36t
5t+36=48-------------------shud be 5t+36t=48
41t=48
41t/41=48/41
t= 48/41
t=1.17
B: Prove your solution is correct?
thank you :)
OK
5*(1.17)/12=4-3*(1.17)
0.4875=4-3.51
0.4875~~~~0.49
Hope this helps---ptaylor